# Simulation-based inference using surjective sequential neural likelihood estimation

### v2

@@ D.4 Estimation of divergences
Zhao et al. (2022) propose using the H-divergence
@@ D.5 Estimation of intrinsic dimensionality
We computed the intrinsic dimensionality of a data set using the "tight local intrinsic dimensionality estimator" (TLE) algorithm (Amsaleg et al., 2019, 2022). The TLE is an estimator of the local intrinsic dimensionality, i.e., the intrinsic dimension of each data point in a data set. Mathematically, the local intrinsic dimensionality for a data point $x$ w.r.t to the distance $r:=r(x)$ to its $k$ nearest neighbors is defined as
@@
For details and the derivation of the TEL intrinsic dimensionality estimator, we refer the reader to Amsaleg et al. (2022).
For Figure 5c, we simulated $N=$1000$$ observations $\{y_{n}\}_{n=1}^{N}$ from the generative models of the Ornstein-Uhlenbeck, Lotka-Volterra and SIR models, respectively, and estimated the local intrinsic dimensions of data set using the Python package scikit-dimension (Bac et al., 2021). The package contains several different estimators for local intrinsic dimensionality, and we chose the TLE estimator arbitrarily.
@@ D.6 Source code
Source code including detailed instructions to reproduce and replicate all experiments can be found in the supplemental material or on GitHub (after publication).
@@ E.1 Simple likelihood complex posterior
The simple likelihood complex posterior (SLCP, Papamakarios et al. (2019)) model with $8$ dimensions uses the following generative process
@@ E.2 Ornstein-Uhlenbeck
The Ornstein-Uhlenbeck (OU) process (Särkkä & Solin, 2019) is a one-dimensional stochastic differential equation that models velocity of a particle suspended in a medium. It has the following form:
@@ E.3 Lotka-Volterra
We follow the parameterization in Lueckmann et al. (2021), but sample a longer time series, i.e., $50$ equally-spaced observations $Y_{t}$ where $t\in\{0,\dots,30\}$. We then concatenate the two $50$-dimensional vectors $y_{t}=[y_{t1}^{T},y_{t2}^{T}]^{T}$ yielding a $100$-dimensional observation.
@@ E.4 SIR model
The SIR model is a model from epidemiology that describes the dynamics of the number of individuals in three compartmental states (susceptible, infectious, or recovered) which is, for instance, be aplpied to model the spread of diseases. We again adopt the presentation by Lueckmann et al. (2021) which defines the generative model
@@
Previous work, e.g., Lueckmann et al. (2021), used continuous normalizing flows (i.e., pushforwards from a continuous base distribution, not continuous normalizing flows as in Chen et al. (2018); Grathwohl et al. (2019)). Continuous likelihood-based models, such as SNL using MAFs, cannot adequately represent discrete data. As a remedy, we dequantize the counts uniformly after sampling them (Theis et al., 2016), i.e., we add noise $u_{t}\sim\mathcal{U}(0,1)$, such that $\tilde{y}_{t}=y_{t}+u_{t}$ and use the noised data for trained. While other approaches to dequantization, such as Ho et al. (2019), would possibly be more rigorous, we found that this simple approach suffices.
@@ E.5 Beta generalized linear model
where $c\in\mathbb{R}^{+}$ is a non-negative concentration parameter, $B$ is computed as in Appendix T.6 of Lueckmann et al. (2021), Beta describes a Beta distribution with a mean-concentration parameterization (Ferrari & Cribari-Neto, 2004)
In Lueckmann et al. (2017), a Bernoulli GLM is used instead of a Beta GLM(called Bernoulli GLM raw). We changed the Bernoulli likelihood to a Beta likelihood because of the the same reason as for the Lotka-Volterra model (a CNF can generally not model a discrete distribution). We followed the implemtation in SBIBM (https://github.com/sbi-benchmark/sbibm) as a design matrix $B$,
@@ E.6 Gaussian mixture model
where $\sigma^{2}=0.01$ $I$ is a unit matrix, and both $\theta\in\mathbb{R}^{2}$ and $Y\in\mathbb{R}^{2}$ are two-dimensional random variables. The GGM again follows the representation in Lueckmann et al. (2021).
@@ E.7 Hyperboloid
The hyperboloid model (Forbes et al., 2022) described in "Negative examples and limitations" in Section 4 is a 2-component mixture model of $t$-distributions of the form
@@
where $t$ represents a Student’s $t$-distribution with $\nu$ degrees of freedom, mean $F(\theta;x)=\left(||\theta-x_{1}||_{2}-||\theta-x_{2}||_{2}\right)$ and scale matrix $\sigma^{2}I$ and $\mathbb{I}$ is vector of ones. We follow Forbes et al. (2022), and in our experiments set $\theta\in\mathbb{R}^{2}$ to be uniformly distribution, $a_{1}=[-0.5,0.0]^{T}$, $a_{2}=[0.5,0.0]^{T}$, $b_{1}=[0.0,-0.5]^{T}$, $a_{2}=[0.0,0.5]^{T}$, $\nu=3$ and $\sigma^{2}=0.01$.
@@ E.8 Solar dynamo
where $\text{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}\exp(-t^{2})\mathrm{d}y$ is the Gauss error function. We simulate a time series of length $N=100$ recursively starting from $y_{0}=[1,1]^{T}$. We follow Albert et al. (2022), and set $b_{1}=0.6$, $w_{1}=0.2$, $b_{2}=1$ and $w_{2}=0.8$.
@@ E.9 Neural mass model
The stochastic versionJansen-Rit neural mass model (Buckwar et al., 2020) describes the collective electrical activity of a neurons. The model is a $6$-dimensional stochastic differential equation of the form
@@
$\theta_{3}$ $\sim\mathcal{U}(0,500)$
$\theta_{4}$ $\sim\mathcal{U}(-30,30)$
@@
We are interest in inference of the $4$-dimensional vector $\theta=[\theta_{1},\dots,\theta_{4}]^{T}=[C,\mu,\sigma,g]^{T}$. The parameters $C_{i}$ are related via $C_{1}=\theta_{1}$, $C_{2}=0.8\theta_{1}$, $C_{3}=0.25$ and $C_{4}=0.25\theta_{1}$. The other parameters are $\mu_{4}=\theta_{2}$, $\sigma_{4}=\theta_{3}$ and $g=\theta_{4}$. We initialize $y_{0}=[0,\dots,0]^{T}$ and simulate 100 equally-spaced observations $Y_{t}$ with $t\in[0,1]$.
We refer the reader to Buckwar et al. (2020) and Ableidinger et al. (2017) for detailed explanations of all constants and equations from where we also adopted the parameterization: $A=3.25$, $B=22$, $a=100$, $b=100$, $v_{\text{max}}=5$, $v_{0}=6$, $r=0.56$ and $\sigma_{5}=\sigma_{6}=2000$ (see also Linhart et al. (2023)).
We solve the SDE using the Python library Diffrax, using an Euler solver.
@@ F.1 Four SBI models benchmark
(a) H-Min divergences
(b) H-Jensen Shannon divergences.
Figure 9: H-Min and H-Jenson Shannon divergences of the Ornstein-Uhlenbeck, Lotka-Volterra, SIR and Beta GLM models. SSNL consistently outperforms SNL, SNRE-C and SNPE-C on Ornstein-Uhlenbeck, Lotka-Volterra, is on par with SNL on Beta GLM and displays mixed results on SIR on both divergence measure. Given that SSNL requires less parameters than SNL, SSNL has the clear advantage in Ornstein-Uhlenbeck, Lotka-Volterra and Beta GLM. The two divergences are consistent for three of the four models, for SIR the H-Min and H-Jensen Shannon show inconsistent divergences.
@@ F.2 Negative examples
Figure 10: Negative examples. We show the H-Min and H-Jensen Shannon divergences on the Gaussian mixture and hyperboloid models. In both cases, SSNL can not outperform the three baselines. Since all data dimensions are informative of posterior parameters, reducing the dimensionality is theoretically only detrimental to the inferences.
@@ F.3 Solar dynamo model
Figure 13: Solar dynamo model method errors (in larger). The $y$-axis shows the mean squared error of a posterior sample and the prior parameters $\theta_{0}$ that have been used to simulate $y_{0}\leftarrow sim(\theta_{0})$. It is further normalized by the lowest MSE in the visualized dataset for scaling.
Figure 14: Posterior distributions of SSNL and SNL. The figures show the posterior distributions for a random seed after having trained the models 1 time and 15 times. Black dots/lines correspond to true values of the parameter $\theta$. For this experiment, SSNL already after the first round manages to locate the posterior quite well around the true parameter values. After round 15, both methods have similar posterior distributions which impressively highlights the positive effect of sequential training.
@@ F.4 Jansen-Rit neural mass model
Figure 16: Likelihood profiles for the Jansen-Rit model with different reduction factors. Each profile corresponds to the likelihood on the validation set used during training. The likelihood profiles all show very similar losses (see $y$-axis). As a consequence, we used the model with the greatest reduction on dimensionality, i.e., $25\%$ which reduces the dimensionality in the embedding layer to $Q=25$.
(a) Mean statistic
(b) Median statistic
Figure 17: Jansen-Rit neural mass model method errors (in larger). The $y$-axis shows the mean squared error of a posterior sample and the prior parameters $\theta_{0}$ that have been used to simulate $y_{0}\leftarrow sim(\theta_{0})$. It is further normalized by the lowest MSE in the visualized dataset for scaling. (a) Dots show mean statistics and bars its standard errors. (b) Dots represent median statistics and bars its standard errors.
(a) SSNL posterior distributions.
(b) SNPE-C posterior distributions.
Figure 18: Posterior distributions of SSNL and SNPE-C. The figures show the posterior distributions for a random seed after having trained the models 1 time and 15 times. Black dots/lines correspond to true values of the parameter $\theta$. After the first round both methods fail to correctly infer the true parameter values. After 15 rounds, SSNL gets closer to the truth but, due to the sequential training process, infers multi-modal posteriors. SNPE-C, on the other hand, has significantly broader posterior quantiles except for $\theta_{4}$ which it adequately infers.

### v3

@@ D.4 Estimation of divergences
Zhao et al., (2022) propose using the H-divergence
@@ D.5 Estimation of intrinsic dimensionality
We computed the intrinsic dimensionality of a data set using the "tight local intrinsic dimensionality estimator" (TLE) algorithm (Amsaleg et al.,, 2019, 2022). The TLE is an estimator of the local intrinsic dimensionality, i.e., the intrinsic dimension of each data point in a data set. Mathematically, the local intrinsic dimensionality for a data point $x$ w.r.t to the distance $r:=r(x)$ to its $k$ nearest neighbors is defined as
@@
For details and the derivation of the TEL intrinsic dimensionality estimator, we refer the reader to Amsaleg et al., (2022).
For Figure 5c, we simulated $N=$1000$$ observations $\{y_{n}\}_{n=1}^{N}$ from the generative models of the Ornstein-Uhlenbeck, Lotka-Volterra and SIR models, respectively, and estimated the local intrinsic dimensions of data set using the Python package scikit-dimension (Bac et al.,, 2021). The package contains several different estimators for local intrinsic dimensionality, and we chose the TLE estimator arbitrarily.
@@ D.6 Source code
Source code including detailed instructions to reproduce and replicate all experiments can be found in the supplemental material or on GitHub at github.com/dirmeier/ssnl.
@@ E.1 Simple likelihood complex posterior
The simple likelihood complex posterior (SLCP, Papamakarios et al., (2019)) model with $8$ dimensions uses the following generative process
@@ E.2 Ornstein-Uhlenbeck
The Ornstein-Uhlenbeck (OU) process (Särkkä and Solin,, 2019) is a one-dimensional stochastic differential equation that models velocity of a particle suspended in a medium. It has the following form:
@@ E.3 Lotka-Volterra
We follow the parameterization in Lueckmann et al., (2021), but sample a longer time series, i.e., $50$ equally-spaced observations $Y_{t}$ where $t\in[0,30]$. We then concatenate the two $50$-dimensional vectors $y_{t}=[y_{t1}^{T},y_{t2}^{T}]^{T}$ yielding a $100$-dimensional observation.
@@ E.4 SIR model
The SIR model is a model from epidemiology that describes the dynamics of the number of individuals in three compartmental states (susceptible, infectious, or recovered) which is, for instance, be aplpied to model the spread of diseases. We again adopt the presentation by Lueckmann et al., (2021) which defines the generative model
@@
Previous work, e.g., Lueckmann et al., (2021), used continuous normalizing flows (i.e., pushforwards from a continuous base distribution, not continuous normalizing flows as in Chen et al., (2018); Grathwohl et al., (2019)). Continuous likelihood-based models, such as SNL using MAFs, cannot adequately represent discrete data. As a remedy, we dequantize the counts uniformly after sampling them (Theis et al.,, 2016), i.e., we add noise $u_{t}\sim\mathcal{U}(0,1)$, such that $\tilde{y}_{t}=y_{t}+u_{t}$ and use the noised data for trained. While other approaches to dequantization, such as Ho et al., (2019), would possibly be more rigorous, we found that this simple approach suffices.
@@ E.5 Beta generalized linear model
where $c\in\mathbb{R}^{+}$ is a non-negative concentration parameter, $B$ is computed as in Appendix T.6 of Lueckmann et al., (2021), Beta describes a Beta distribution with a mean-concentration parameterization (Ferrari and Cribari-Neto,, 2004)
In Lueckmann et al., (2017), a Bernoulli GLM is used instead of a Beta GLM(called Bernoulli GLM raw). We changed the Bernoulli likelihood to a Beta likelihood because of the the same reason as for the Lotka-Volterra model (a CNF can generally not model a discrete distribution). We followed the implemtation in SBIBM (https://github.com/sbi-benchmark/sbibm) as a design matrix $B$.
@@ E.6 Gaussian mixture model
where $\sigma^{2}=0.01$ $I$ is a unit matrix, and both $\theta\in\mathbb{R}^{2}$ and $Y\in\mathbb{R}^{2}$ are two-dimensional random variables. The GGM again follows the representation in Lueckmann et al., (2021).
@@ E.7 Hyperboloid
The hyperboloid model (Forbes et al.,, 2022) described in "Negative examples and limitations" in Section 4 is a 2-component mixture model of $t$-distributions of the form
@@
where $t$ represents a Student’s $t$-distribution with $\nu$ degrees of freedom, mean $F(\theta;x)=\left(||\theta-x_{1}||_{2}-||\theta-x_{2}||_{2}\right)$ and scale matrix $\sigma^{2}I$ and $\mathbb{I}$ is vector of ones. We follow Forbes et al., (2022), and in our experiments set $\theta\in\mathbb{R}^{2}$ to be uniformly distribution, $a_{1}=[-0.5,0.0]^{T}$, $a_{2}=[0.5,0.0]^{T}$, $b_{1}=[0.0,-0.5]^{T}$, $a_{2}=[0.0,0.5]^{T}$, $\nu=3$ and $\sigma^{2}=0.01$.
@@ E.8 Solar dynamo
where $\text{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}\exp(-t^{2})\mathrm{d}y$ is the Gauss error function. We simulate a time series of length $N=100$ recursively starting from $y_{0}=[1,1]^{T}$. We follow Albert et al., (2022), and set $b_{1}=0.6$, $w_{1}=0.2$, $b_{2}=1$ and $w_{2}=0.8$.
@@ E.9 Neural mass model
The stochastic version of the Jansen-Rit neural mass model (Ableidinger et al.,, 2017) describes the collective electrical activity of neurons. The model is a $6$-dimensional stochastic differential equation of the form
@@
$\theta_{3}$ $\sim\mathcal{U}(100,5000)$
$\theta_{4}$ $\sim\mathcal{U}(-20,20)$
@@
We are interest in inference of the $4$-dimensional vector $\theta=[\theta_{1},\dots,\theta_{4}]^{T}=[C,\mu,\sigma,g]^{T}$. The parameters $C_{i}$ are related via $C_{1}=\theta_{1}$, $C_{2}=0.8\theta_{1}$, $C_{3}=0.25$ and $C_{4}=0.25\theta_{1}$. The other parameters are $\mu_{4}=\theta_{2}$, $\sigma_{5}=\theta_{3}$ and $g=\theta_{4}$. Following previous work, we initialize $y_{0}=[0.08,18,15,-0.5,0,0]^{T}$ and simulate a time series $Y_{t}$ with $t\in[0,8]$ with sampling frequency $Hz=512$. We then takes $100$ equally-spaced elements from $Y_{t}$.
We refer the reader to Ableidinger et al., (2017), Rodrigues et al., (2021) and Buckwar et al., (2020) for detailed explanations of all constants and equations from where we also adopted the parameterization: $A=3.25$, $B=22$, $a=100$, $b=50$, $v_{\text{max}}=5$, $v_{0}=6$, $r=0.56$, $\sigma_{4}=0.01$ and $\sigma_{6}=1$ (see also Linhart et al., (2023)).
We solve the SDE with the Python library jrnmm using the Strang-splitting method as described in Buckwar et al., (2020).
@@ F.1 Four SBI models benchmark
(a) H-Min divergences.
(b) H-Min divergences (all baselines).
(c) H-Jensen Shannon divergences.
(d) H-Jensen Shannon divergences (all baselines).
Figure 9: H-Min and H-Jenson Shannon divergences of the Ornstein-Uhlenbeck, Lotka-Volterra, SIR and Beta GLM models (left withouth SNPE-C for Lotka-Volterra and SIR, right with all baselines). SSNL consistently outperforms all five baselines on Ornstein-Uhlenbeck, Lotka-Volterra, is on par with SNL on Beta GLM and displays mixed results on SIR on both divergence measure. Given that SSNL requires less parameters than SNL, SSNL has the clear advantage in Ornstein-Uhlenbeck, Lotka-Volterra and Beta GLM. The two divergences are consistent for three of the four models, for SIR the H-Min and H-Jensen Shannon show inconsistent divergences.
@@ F.2 Negative examples
Figure 10: Negative examples. We show the H-Min and H-Jensen Shannon divergences on the Gaussian mixture and hyperboloid models. In both cases, SSNL can not outperform the three baselines. Since all data dimensions are informative of posterior parameters, reducing the dimensionality is theoretically only detrimental to the inferences. We did not conduct experiments on SNASS and SNASSS here, since their is no dimensionality reduction necessary.
@@ F.4 Jansen-Rit neural mass model
Figure 14: Likelihood profiles for the Jansen-Rit model with different reduction factors. Each profile corresponds to the likelihood on the validation set used during training. We used the model with the greatest reduction on dimensionality, i.e., $25\%$, which reduces the dimensionality in the embedding layer to $Q=25$.